# Topological zero modes and bounded modes at smooth domain walls: exact solutions and dualities

### v1

@@ Introduction
$\left(\left(\eta p^{2}+m(x)\right)\sigma_{z}+{2v}(x)p\,\sigma_{y}\right)\Psi(x)=E\Psi(x).$ (1)
where $\bm{\Psi}(x)^{\dagger}=[\Psi(x)^{\dagger},\Psi(x)]$ is a spinor with $\Psi(x)$ a fermion field, $\sigma_{xyz}$ the Pauli matrices, $p=-\mathrm{i}\partial$ the momentum operator, $\eta\geq 0$, and with $m(x),v(x)\in\mathbb{R}$. This equation describes a Dirac fermion of mass $m(x)$ and speed of light $2v(x)$ if $\eta=0$. In condensed matter, this equation describes a topological insulator with effective mass $1/2\eta$, chemical potential $\mu(x)=-m(x)$, and spin-orbit coupling $2v(x)$, or a spinless topological superconductor with effective mass $1/2\eta$, chemical potential $\mu(x)=-m(x)$, and $p$-wave superconducting pairing $\Delta_{p}(x)=2v(x)$. Here, we only focus on the zero modes $E=0$ which are normalizable and satisfy Dirichlet boundary conditions on infinite or semi-infinite intervals. These modes are necessarily eigenstates of $\sigma_{x}$ in the form
$\bm{\psi}(x)\propto\begin{bmatrix}1\\ s\end{bmatrix}\varphi(x),\,\,\sigma_{x}\bm{\psi}(x)=s\bm{\psi}(x),$ (2)
@@
where we absorbed $\eta$ into the definition of the fields $v(x)\to v(x)/\eta$ and $m(x)\to m(x)/\eta$. We refer to the eigenvalue $s$ as the pseudospin. Notice that the equation is real with real solutions and consequently the zero modes are Majorana modes analogous to the Majorana solutions of the Dirac equations.
@@
For uniform fields $m(x)=m$, $v(x)=v$, one can define a topological invariant $w\in\mathbb{Z}_{2}$ characterizing the Hamiltonian in Eq. 1, which is $w=\operatorname{sgn}(v)$ for $m<0$ and $w=0$ otherwise[14]. Due to the bulk-boundary correspondence[11, 12, 13], zero energy modes localize at the boundary between topologically trivial $w=0$ and nontrivial regions $w\neq 0$, or at the boundary between topologically nontrivial phases with different topological invariants. The number of zero energy modes localized at the boundary is equal to the difference between the topological invariant on the right and on the left of the boundary $\Delta w=|w_{L}-w_{R}|$.
@@
We assume that the fields approach their limit values on the left and right exponentially in the system size $|m(x\to\mp\infty)-m_{L,R}|\propto\mathrm{e}^{-|x|/w}$, $|v(x\to\mp\infty)-v_{L,R}|\propto\mathrm{e}^{-|x|/w}$, where $w>0$ is a characteristic length describing the spatial variations of fields, i.e., the width of the smooth domain wall localized at the origin $x=0$. By mapping the whole real line $-\infty<x<\infty$ into the finite segment $10<y<$ 1 with the substitution $y(x)=(1+\tanh{(x/2w)})/2$, we solve the equation around the singular points $y=0,1$. We found that for a given pseudospin $s=\pm 1$, the general solution of the modified Jackiw-Rebbi equation is given by a linear combination $\sum_{i}A_{i}\varphi_{i}^{s}(x)$ where
$\varphi_{1,2}^{s}(x)=$ $\frac{\mathrm{e}^{({\alpha_{L}^{\pm}}-{\alpha_{R}^{\mp}})x/2}}{\left(2\cosh{(x/2w)}\right)^{{w(\alpha_{L}^{\pm}}+{\alpha_{R}^{\mp}})}}F_{1,2}(x),$ (4)
@@
with exponents $\alpha_{L}^{\pm}=-s{v}_{L}\pm q_{L}$, $\alpha_{R}^{\pm}=s{v}_{R}\pm q_{R}$, where $q_{L,R}=\sqrt{v_{L,R}^{2}+m_{L,R}}$, which depend only on the values of the fields at large distances $|x|\gg w$, and where the functions $F_{1,2}$ are bounded or diverge polynomially for $x\to\pm\infty$. Asymptotically, these solutions give $\varphi_{1,2}^{s}(x\to-\infty)\propto\mathrm{e}^{-\alpha_{L}^{\pm}|x|}=\mathrm{e}^{-\mu_{L}^{\pm}|x|}\mathrm{e}^{\mp\mathrm{i}\kappa_{L}|x|}$, $\varphi_{1,2}^{s}(x\to+\infty)\propto\mathrm{e}^{-\alpha_{R}^{\mp}|x|}=\mathrm{e}^{-\mu_{R}^{\mp}|x|}\mathrm{e}^{\pm\mathrm{i}\kappa_{R}|x|}$, where $\mu_{L,R}^{\pm}=1/\xi_{L,R}^{\pm}=\mathrm{Re}(\alpha_{L,R}^{\pm})$ are the decay rates and decay lengths, while $\kappa_{L,R}=2\pi/\lambda_{L,R}=\mathrm{Im}(q_{L,R})$ the momentum and wavelengths responsible for oscillatory behaviors. One has $\kappa_{L,R}=0$ for ${v}_{L,R}^{2}+m_{L,R}\geq 0$, with $\mu_{L,R}^{+}=\mu_{L,R}^{-}$ for ${v}_{L,R}^{2}+m_{L,R}=0$, while $\mu_{L,R}^{+}=\mu_{L,R}^{-}=\mp sv_{L,R}$ and $\kappa_{L,R}=\mathrm{Im}(q_{L,R})\neq 0$ for ${v}_{L,R}^{2}+m_{L,R}<0$.
@@
In the case where the fields $v(x)$ and $m(x)$ can be expanded in powers of $y(x)=\left(1+\tanh{(x/2w)}\right)/2$ up to the first and second order, respectively, i.e., as $v(x)=v_{0}+v_{1}y(x)$, and $m(x)=m_{0}+m_{1}y(x)+m_{2}y(x)^{2}$, where $v_{0}=v_{L}$, $v_{1}=v_{R}-v_{L}$, $m_{0}=m_{L}$, $m_{1}+m_{2}=m_{R}-m_{L}$, and $m_{2}=2(m_{L}+m_{R})-4m_{D}$, with $m_{D}=m(0)$, we can transform the equation into a well-known hypergeometric equation, and solve it in terms of hypergeometric functions. Indeed, for a given pseudospin $s=\pm 1$, we found that
@@
$a_{1,2}$ =±$=\pm w\left({q_{L}}-{q_{R}}\right)+\tfrac{1}{2}+\tfrac{1}{2}{\sqrt{\left(2wsv_{1}-1\right)^{2}+4w^{2}m_{2}}}$ (6a)
$b_{1,2}$ =±$=\pm w\left({q_{L}}-{q_{R}}\right)+\tfrac{1}{2}-\tfrac{1}{2}{\sqrt{\left(2wsv_{1}-1\right)^{2}+4w^{2}m_{2}}}$ (6b)
@@
The solutions $\varphi_{1,2}^{s}(x)$ simplify in the cases where $|v_{L,R}|=v$ or $m_{L,R}=m$ or if the fields are uniform, as shown in Table 1. The functions $F_{1,2}$ are bounded or diverge polynomially for all choices of the parameters. Hence, the exponents $\alpha_{L}^{\pm}$ and $\alpha_{R}^{\pm}$ uniquely determine the asymptotic behavior of the solutions and whether there exist particular solutions that satisfy the boundary conditions on a given interval, i.e., whether zero modes exist on a finite, semi-infinite, or infinite system. These exponents depend uniquely on the values of the fields at large distances $m_{L,R}$ and $v_{L,R}$. The parameters $a_{1,2}$, $b_{1,2}$, and $c_{1,2}$ depend as well on the values of the fields at large distances $m_{L,R}$ and $v_{L,R}$. However, the two parameters $a_{1,2}$ and $b_{1,2}$ depend also on he values of the field $m(x)$ at the domain wall via $m_{2}$, since $m_{2}=2(m_{L}+m_{R})-4m_{D}$ where $m_{D}=m(x=0)$ is the value of the field $m(x)$ at the center of the domain wall $x=0$.
1) general case $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\left({\alpha_{L}^{\pm}}-{\alpha_{R}^{\mp}}\right)x/2}}{\left(2\cosh{(x/2w)}\right)^{w\left(\alpha_{L}^{\pm}+\alpha_{R}^{\mp}\right)}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$=\dfrac{\mathrm{e}^{\left[\pm\left(q_{L}+q_{R}\right)-s\left(v_{L}+v_{R}\right)\right]\,x/2}}{\left(2\cosh{(x/2w)}\right)^{w\left[\pm\left(q_{L}-{q_{R}}\right)+sv_{1}\right]}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
2) $v(x)=v$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{-sv\,x}\mathrm{e}^{\pm\left(q_{L}+q_{R}\right)x/2}}{\left(2\cosh{(x/2w)}\right)^{\pm w\left(q_{L}-q_{R}\right)}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
5) $v_{L,R}=\pm v$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\pm\left(q_{L}+q_{R}\right)x/2}}{\left(2\cosh{(x/2w)}\right)^{\pm w\left(q_{L}-q_{R}\right)-2wsv}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
6) $v_{L,R}=\pm v$, $m_{L,R}=m$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\pm\sqrt{{v}^{2}+m}\,\,x}}{\left(2\cosh{(x/2w)}\right)^{-2wsv}}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
7) $v_{L,R}=\pm v$, $m(x)=m$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\pm\sqrt{{v}^{2}+m}\,\,x}}{\left(2\cosh{(x/2w)}\right)^{-2wsv}}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
8) $q_{L,R}=q$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\left[\pm 2q-s(v_{L}+v_{R})\right]\,x/2}}{\left(2\cosh{(x/2w)}\right)^{wsv_{1}}}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
The values of the fields at large distances and the boundary conditions univocally determine the existence and number of particular solutions and their pseudospin $s$ on infinite and semi-infinite intervals. In particular, to ascertain the existence, number, and asymptotic properties of the solutions we define the topological invariants
@@
These topological invariants depend only on the values of the fields $m_{L,R}$ and $v_{L,R}$ at large distances $|x|>w$. The phase on the left for $x\ll-w$ is topologically nontrivial if $\operatorname{sgn}(m_{L})=-1$ with topological invariant equal to $W_{L}=\operatorname{sgn}(v_{L})$ and trivial otherwise, while the phase on the right of the smooth domain wall for $x\gg w$ is topologically nontrivial if $\operatorname{sgn}(m_{R})=-1$ with topological invariant equal to $W_{R}=\operatorname{sgn}(v_{R})$ and trivial otherwise (see Refs. 14, 8).
Zero modes on the interval $-\infty<x<\infty$ satisfying the boundary conditions $\varphi(\pm\infty)=0$ are given by the modes $\varphi_{1,2}^{s}(x)$ in Eq. 4 with $\mathrm{Re}(\alpha_{L}^{\pm})>0$ and $\mathrm{Re}(\alpha_{R}^{\pm})>0$. As one would expect from the bulk-boundary correspondence[11, 12, 13], there are two independent modes if $W_{L}W_{R}=-1$, one independent mode if $W_{L}=\pm 1$ and $W_{R}=0$ or if $W_{R}=\pm 1$ and $W_{L}=0$.
For $\operatorname{sgn}(m_{L,R})=-1$ and $W_{L}W_{R}=-1$, taking $s=-W_{L}=W_{R}$ gives $\mathrm{Re}(\alpha_{L}^{\pm}),\mathrm{Re}(\alpha_{R}^{\pm})>0$: Hence, both $\varphi_{1,2}^{s}(x)$ decays exponentially to zero for $x\to\pm\infty$, and consequently the particular solution satisfying the boundary conditions is any linear combination $\varphi(x)=\sum_{i}A_{i}\varphi_{i}^{s}(x)$. The asymptotic behavior of the modes $\varphi_{1}^{s}$ and $\varphi_{2}^{s}$ are determined on the left side by the exponents $\alpha_{L}^{+}=-sv_{L}+q_{L}$ (characteristic lengths $\xi_{L}^{+}$, $\lambda_{L}^{+}$) and $\alpha_{L}^{-}=-sv_{L}-q_{L}$ (characteristic lengths $\xi_{L}^{-}$, $\lambda_{L}^{-}$), respectively, and on the right side by the exponents $\alpha_{R}^{-}=sv_{R}-q_{R}$ (characteristic lengths $\xi_{R}^{-}$, $\lambda_{R}^{-}$) and $\alpha_{R}^{+}=sv_{R}+q_{R}$ (characteristic lengths $\xi_{R}^{+}$, $\lambda_{R}^{+}$), respectively.
@@
An analogous argument holds for $\operatorname{sgn}(m_{R})=1$, $\operatorname{sgn}(m_{L})=-1$ (i.e., $W_{R}=0$ and $|W_{L}|=1$), and for the limiting case $\operatorname{sgn}(m_{R})=0$, $\operatorname{sgn}(m_{L})=-1$ with $W_{L}W_{R}=-1$, taking $s=-W_{L}$, giving a single mode satisfying the boundary conditions $\varphi(x)=A_{2}\varphi_{2}^{s}(x)$, with asymptotic behavior determined on the left by the exponents $\alpha_{L}^{-}=-sv_{L}-q_{L}$ (characteristic lengths $\xi_{L}^{-}$, $\lambda_{L}^{-}$) and on the right by $\alpha_{R}^{+}=sv_{R}+q_{R}$ (characteristic lengths $\xi_{R}^{+}$, $\lambda_{R}^{+}$).
Zero modes on the interval $0\leq x<\infty$ satisfying the boundary conditions $\varphi(0)=\varphi(\infty)=0$ are given by linear combinations of $\varphi_{1,2}^{s}(x)$ in Eq. 4 with $\mathrm{Re}(\alpha_{R}^{\pm})>0$. As one would expect from the bulk-boundary correspondence[11, 12, 13], there is one independent mode if $W_{R}\neq 1$. For $\operatorname{sgn}(m_{R})=-1$ and $|W_{R}|=1$, choosing $s=W_{R}$ gives $\mathrm{Re}(\alpha_{R}^{\pm})>0$, and therefore both $\varphi_{1,2}^{s}(x)$ decays to zero for $x\to\infty$, and thus one can obtain a single particular solution satisfying the boundary conditions, which is the linear combination $\varphi(x)=\sum_{i}A_{i}\varphi_{i}^{s}(x)$ with $A_{1,2}$ chosen such that $\varphi(0)=0$. The asymptotic behavior is determined by the exponent $\alpha_{R}^{-}=sv_{R}-q_{R}$ (characteristic lengths $\xi_{R}^{-}$, $\lambda_{R}^{-}$). This is because, even thought this mode is a linear combination of two exponentially decaying solutions with decaying lengths $\xi_{R}^{\pm}$, one has that $\xi_{R}^{-}\geq\xi_{R}^{+}$, which means that the mode $\varphi_{1}^{s}$ with decaying length $\xi_{R}^{-}$ dominates at large distances. In the special cases where $|v_{L,R}|=v$ and $m_{R}=m_{L}$, one has ${A_{1}}/{A_{2}}=-1$. Zero modes on the interval $-\infty<x\leq 0$ can be obtained similarly.

### v2

@@ Introduction
$\left(\left(\eta p^{2}+m(x)\right)\sigma_{z}+{2v}(x)p\,\sigma_{y}\right)\Psi(x)=E\Psi(x),$ (1)
where $\bm{\Psi}(x)^{\dagger}=(\Psi(x)^{\dagger},\Psi(x))$ is a spinor with $\Psi(x)$ a fermion field, $\sigma_{xyz}$ the Pauli matrices, $p=-\mathrm{i}\partial$ the momentum operator, $\eta\geq 0$, and with $m(x),v(x)\in\mathbb{R}$. This equation describes a Dirac fermion of mass $m(x)$ and speed of light $2v(x)$ if $\eta=0$. In condensed matter, this equation describes a topological insulator with effective mass $1/2\eta$, chemical potential $\mu(x)=-m(x)$, and spin-orbit coupling $2v(x)$, or a spinless topological superconductor with effective mass $1/2\eta$, chemical potential $\mu(x)=-m(x)$, and $p$-wave superconducting pairing $\Delta_{p}(x)=2v(x)$. Here, we only focus on the zero modes $E=0$ which are normalizable and satisfy Dirichlet boundary conditions on infinite or semi-infinite intervals. These modes are necessarily eigenstates of $\sigma_{x}$ in the form
$\bm{\psi}(x)\propto\begin{pmatrix}1\\ s\end{pmatrix}\varphi(x),\,\,\sigma_{x}\bm{\psi}(x)=s\bm{\psi}(x),$ (2)
@@
where we absorbed $\eta$ into the definition of the fields $v(x)\to v(x)/\eta$ and $m(x)\to m(x)/\eta$. We refer to the eigenvalue $s$ as the pseudospin. Notice that the equation is real with real solutions and consequently the zero modes are Majorana modes analogous to the Majorana solutions of the Dirac equations. Hence, the real eigenstates of the equation above are also eigenstates of the antiunitary complex conjugation operator $K$ and of the antiunitary operator $C=K\sigma_{x}$, which can be interpreted as the antiunitary time-reversal and charge conjugation (or particle-hole symmetry in condensed matter).
@@
For uniform fields $m(x)=m$, $v(x)=v$, one can define a topological invariant $W\in\mathbb{Z}_{2}$ characterizing the Hamiltonian in Eq. 1, which is $W=\operatorname{sgn}(v)$ for $m\leq 0$ and $W=0$ otherwise[14]. Due to the bulk-boundary correspondence[11, 12, 13], zero energy modes localize at the boundary between topologically trivial $W=0$ and nontrivial regions $W\neq 0$, or at the boundary between topologically nontrivial phases with different topological invariants. The number of zero energy modes localized at the boundary is equal to the difference between the topological invariant on the right and on the left of the boundary $\Delta W=|W_{L}-W_{R}|$.
@@
We assume that the fields approach their limit values on the left and right exponentially in the system size $|m(x\to\mp\infty)-m_{L,R}|\propto\mathrm{e}^{-|x|/w}$, $|v(x\to\mp\infty)-v_{L,R}|\propto\mathrm{e}^{-|x|/w}$, where $w>0$ is a characteristic length describing the spatial variations of fields, i.e., the width of the smooth domain wall localized at the origin $x=0$. By mapping the whole real line $-\infty<x<\infty$ into the finite segment $10<y<$ 1 with the substitution $y(x)=(1+\tanh{(x/2w)})/2$, one obtains
$\varphi^{\prime\prime}(y)+\frac{1-2y+2wsv(y)}{y(1-y)}\varphi^{\prime}(y)-\frac{w^{2}m(y)}{y^{2}(1-y)^{2}}\varphi(y)=0,$ (4)
which presents two regular singular points at $y=0,1$. Hence, assuming a solution $\varphi(y)=y^{w\alpha_{L}}(1-y)^{w{\alpha_{R}}}F(y)$ by solving indicial equations for the exponents at the singular points we find the exponents $\alpha_{L}^{\pm}=-s{v}_{L}\pm q_{L}$, $\alpha_{R}^{\pm}=s{v}_{R}\pm q_{R}$, where $q_{L,R}=\sqrt{v_{L,R}^{2}+m_{L,R}}$. Hence, for a given pseudospin $s=\pm 1$, the general solution of the modified Jackiw-Rebbi equation is given by a linear combination $\sum_{i}A_{i}\varphi_{i}^{s}(x)$ where
$\varphi_{1,2}^{s}(x)=$ $\frac{\mathrm{e}^{({\alpha_{L}^{\pm}}-{\alpha_{R}^{\mp}})x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{{w(\alpha_{L}^{\pm}}+{\alpha_{R}^{\mp}})}}F_{1,2}(x),$ (5)
@@
which depend only on the values of the fields at large distances $|x|\gg w$, and where the functions $F_{1,2}$ are bounded or diverge polynomially for $x\to\pm\infty$. Asymptotically, these solutions give $\varphi_{1,2}^{s}(x\to-\infty)\propto\mathrm{e}^{-\alpha_{L}^{\pm}|x|}=\mathrm{e}^{-\mu_{L}^{\pm}|x|}\mathrm{e}^{\mp\mathrm{i}\kappa_{L}|x|}$, $\varphi_{1,2}^{s}(x\to+\infty)\propto\mathrm{e}^{-\alpha_{R}^{\mp}|x|}=\mathrm{e}^{-\mu_{R}^{\mp}|x|}\mathrm{e}^{\pm\mathrm{i}\kappa_{R}|x|}$, where $\mu_{L,R}^{\pm}=1/\xi_{L,R}^{\pm}=\mathrm{Re}(\alpha_{L,R}^{\pm})$ are the decay rates and decay lengths, while $\kappa_{L,R}=2\pi/\lambda_{L,R}=\mathrm{Im}(q_{L,R})$ the momentum and wavelengths responsible for oscillatory behaviors. For ${v}_{L,R}^{2}+m_{L,R}\geq 0$ one has $\kappa_{L,R}=0$, with $\mu_{L,R}^{+}=\mu_{L,R}^{-}$ in the limiting case ${v}_{L,R}^{2}+m_{L,R}=0$. Conversely, for ${v}_{L,R}^{2}+m_{L,R}<0$ one has that $\mu_{L}^{+}=\mu_{L}^{-}=-sv_{L}$, $\mu_{R}^{+}=\mu_{R}^{-}=sv_{R}$, and $\kappa_{L,R}=\mathrm{Im}(q_{L,R})\neq 0$.
@@
In the case where $v(x)$ and $m(x)$ can be expanded in powers of $y(x)=\left(1+\tanh{(x/2w)}\right)/2$ up to the first and second order, respectively, i.e., as $v(x)=v_{0}+v_{1}y(x)$, and $m(x)=m_{0}+m_{1}y(x)+m_{2}y(x)^{2}$, where $v_{0}=v_{L}$, $v_{1}=v_{R}-v_{L}$, $m_{0}=m_{L}$, $m_{1}+m_{2}=m_{R}-m_{L}$, and $m_{2}=2(m_{L}+m_{R})-4m_{D}$, with $m_{D}=m(0)$, we find an explicit general solution in terms of hypergeometric functions. For each set of exponents $({\alpha_{L}},{\alpha_{R}})=(\alpha_{L}^{\pm},\alpha_{R}^{\pm})$ the equation becomes
$F^{\prime\prime}(y)+\frac{c-(a+b+1)y}{y(1-y)}F^{\prime}(y)-\frac{ab}{y(1-y)}F(y)=0,$ (6)
which is the hypergeometric equation with $a,b=w{\alpha_{L}}+w{\alpha_{R}}-wsv_{1}+\frac{1}{2}\pm\frac{1}{2}\sqrt{\left(2wsv_{1}-1\right)^{2}+4w^{2}m_{2}}$ and $c=1+2wsv_{L}+2w{\alpha_{L}}$, The solutions of the hypergeometric equation can be chosen as two linearly independent solutions out of the Kummer’s 24 solutions of the hypergeometric equation, e.g., taking ${\,}_{2}F_{1}\left(a,b,c,y\right)$, and ${\,}_{2}F_{1}\left(a,b,a+b-c+1,1-y\right)$. Hence, for a given pseudospin $s=\pm 1$, we found that
@@
$a_{1,2}$ =±$=\pm w\left({q_{L}}-{q_{R}}\right)+\tfrac{1}{2}+\tfrac{1}{2}{\sqrt{\left(2wsv_{1}-1\right)^{2}+4w^{2}m_{2}}},$ (8a)
$b_{1,2}$ =±$=\pm w\left({q_{L}}-{q_{R}}\right)+\tfrac{1}{2}-\tfrac{1}{2}{\sqrt{\left(2wsv_{1}-1\right)^{2}+4w^{2}m_{2}}},$ (8b)
@@
The solutions $\varphi_{1,2}^{s}(x)$ simplify in the cases where $|v_{L,R}|=v$ or $m_{L,R}=m$ or if the fields are uniform, as shown in Table 1. The functions $F_{1,2}$ are bounded or diverge polynomially for all choices of the parameters. Hence, the exponents $\alpha_{L}^{\pm}$ and $\alpha_{R}^{\pm}$ uniquely determine the asymptotic behavior of the solutions and whether there exist particular solutions that satisfy the boundary conditions on a given interval, i.e., whether zero modes exist on a finite, semi-infinite, or infinite system. These exponents depend uniquely on the values of the fields at large distances $m_{L,R}$ and $v_{L,R}$. The parameters $a_{1,2}$, $b_{1,2}$, and $c_{1,2}$ depend as well on the values of the fields at large distances $m_{L,R}$ and $v_{L,R}$. However, the two parameters $a_{1,2}$ and $b_{1,2}$ also depend on the values of the field $m(x)$ at the domain wall via $m_{2}$, since $m_{2}=2(m_{L}+m_{R})-4m_{D}$ where $m_{D}=m(x=0)$ is the value of the field $m(x)$ at the center of the domain wall $x=0$. Note that the value of $m_{2}$ gives a measure of the variations of the field around the domain wall, since $m_{2}/4$ equals to the difference between the value of the field at $x=0$ and the average of the field at large distances $x\to\pm\infty$.
1) general case $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\left({\alpha_{L}^{\pm}}-{\alpha_{R}^{\mp}}\right)x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{w\left(\alpha_{L}^{\pm}+\alpha_{R}^{\mp}\right)}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$=\dfrac{\mathrm{e}^{\left[\pm\left(q_{L}+q_{R}\right)-s\left(v_{L}+v_{R}\right)\right]\,x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{w\left[\pm\left(q_{L}-{q_{R}}\right)+sv_{1}\right]}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
2) $v(x)=v$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{-sv\,x}\mathrm{e}^{\pm\left(q_{L}+q_{R}\right)x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{\pm w\left(q_{L}-q_{R}\right)}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
5) $v_{L,R}=\pm v$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\pm\left(q_{L}+q_{R}\right)x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{\pm w\left(q_{L}-q_{R}\right)-2wsv}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
6) $v_{L,R}=\pm v$, $m_{L,R}=m$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\pm\sqrt{{v}^{2}+m}\,\,x}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{-2wsv}}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
7) $v_{L,R}=\pm v$, $m(x)=m$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\pm\sqrt{{v}^{2}+m}\,\,x}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{-2wsv}}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
8) $q_{L,R}=q$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\left[\pm 2q-s(v_{L}+v_{R})\right]\,x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{wsv_{1}}}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
@@
The values of the fields at large distances and the boundary conditions univocally determine the existence and number of particular solutions and their pseudospin $s$ on infinite and semi-infinite intervals. In particular, to ascertain the existence, number, and asymptotic properties of the solutions, we define the topological invariants
@@
These topological invariants depend only on the values of the fields $m_{L,R}$ and $v_{L,R}$ at large distances $|x|>w$. The phase on the left $x\ll-w$ is topologically nontrivial if $m_{L}\leq 0$ and $v_{L}\neq 0$ with topological invariant equal to $W_{L}=\operatorname{sgn}(v_{L})$ and trivial otherwise. Analogously, the phase on the right $x\gg w$ is nontrivial if $m_{R}\leq 0$ and $v_{R}\neq 0$ with $W_{R}=\operatorname{sgn}(v_{R})$ and trivial otherwise (see Refs. 14, 8).
Zero modes on the interval $-\infty<x<\infty$ satisfying the boundary conditions $\varphi(\pm\infty)=0$ are given by the modes $\varphi_{1,2}^{s}(x)$ in Eq. 5 with $\mathrm{Re}(\alpha_{L}^{\pm})>0$ and $\mathrm{Re}(\alpha_{R}^{\pm})>0$. As one would expect from the bulk-boundary correspondence[11, 12, 13], there are two independent modes if $W_{L}W_{R}=-1$, one independent mode if $W_{L}=\pm 1$ and $W_{R}=0$ or if $W_{R}=\pm 1$ and $W_{L}=0$. Indeed, for $\operatorname{sgn}(m_{L,R})=-1$ and $W_{L}W_{R}=-1$, taking $s=-W_{L}=W_{R}$ gives $\mathrm{Re}(\alpha_{L}^{\pm}),\mathrm{Re}(\alpha_{R}^{\pm})>0$: Hence, both $\varphi_{1,2}^{s}(x)$ decay exponentially to zero for $x\to\pm\infty$, and consequently the particular solution satisfying the boundary conditions is any linear combination $\varphi(x)=\sum_{i}A_{i}\varphi_{i}^{s}(x)$. The asymptotic behavior of the modes $\varphi_{1}^{s}$ and $\varphi_{2}^{s}$ are determined on the left side by the exponents $\alpha_{L}^{+}=-sv_{L}+q_{L}$ (characteristic lengths $\xi_{L}^{+}$, $\lambda_{L}^{+}$) and $\alpha_{L}^{-}=-sv_{L}-q_{L}$ (characteristic lengths $\xi_{L}^{-}$, $\lambda_{L}^{-}$), respectively, and on the right side by the exponents $\alpha_{R}^{-}=sv_{R}-q_{R}$ (characteristic lengths $\xi_{R}^{-}$, $\lambda_{R}^{-}$) and $\alpha_{R}^{+}=sv_{R}+q_{R}$ (characteristic lengths $\xi_{R}^{+}$, $\lambda_{R}^{+}$), respectively.
@@
Analogous arguments hold for $\operatorname{sgn}(m_{R})=1$, $\operatorname{sgn}(m_{L})=-1$ (i.e., $W_{R}=0$ and $|W_{L}|=1$), and for the limiting case $\operatorname{sgn}(m_{R})=0$, $\operatorname{sgn}(m_{L})=-1$ with $W_{L}W_{R}=-1$, taking $s=-W_{L}$, giving a single mode satisfying the boundary conditions $\varphi(x)=A_{2}\varphi_{2}^{s}(x)$, with asymptotic behavior determined on the left by the exponents $\alpha_{L}^{-}=-sv_{L}-q_{L}$ (characteristic lengths $\xi_{L}^{-}$, $\lambda_{L}^{-}$) and on the right by $\alpha_{R}^{+}=sv_{R}+q_{R}$ (characteristic lengths $\xi_{R}^{+}$, $\lambda_{R}^{+}$).
Zero modes on the interval $0\leq x<\infty$ satisfying the boundary conditions $\varphi(0)=\varphi(\infty)=0$ are given by linear combinations of $\varphi_{1,2}^{s}(x)$ in Eq. 5 with $\mathrm{Re}(\alpha_{R}^{\pm})>0$. As one would expect from the bulk-boundary correspondence[11, 12, 13], there is one independent mode if $W_{R}\neq 1$. For $\operatorname{sgn}(m_{R})=-1$ and $|W_{R}|=1$, choosing $s=W_{R}$ gives $\mathrm{Re}(\alpha_{R}^{\pm})>0$, and therefore both $\varphi_{1,2}^{s}(x)$ decays to zero for $x\to\infty$, and thus one can obtain a single particular solution satisfying the boundary conditions, which is the linear combination $\varphi(x)=\sum_{i}A_{i}\varphi_{i}^{s}(x)$ with $A_{1,2}$ chosen such that $\varphi(0)=0$. The asymptotic behavior is determined by the exponent $\alpha_{R}^{-}=sv_{R}-q_{R}$ (characteristic lengths $\xi_{R}^{-}$, $\lambda_{R}^{-}$). This is because, even though this mode is a linear combination of two exponentially decaying solutions with decaying lengths $\xi_{R}^{\pm}$, one has that $\xi_{R}^{-}\geq\xi_{R}^{+}$, which means that the mode $\varphi_{1}^{s}$ with decaying length $\xi_{R}^{-}$ dominates at large distances. In the special cases where $|v_{L,R}|=v$ and $m_{R}=m_{L}$, one has ${A_{1}}/{A_{2}}=-1$. Zero modes on the interval $-\infty<x\leq 0$ can be obtained similarly.